OpenStudy (micahm):

The data in the table below represents the pressure of the gas as the temperature changes. Plot a graph of this data, using the blank graph provided below. Draw a trend line and calculate its slope. How are the variables related? What will the pressure of the gas be at 0°C?

OpenStudy (micahm):

i found this link http://en.wikibooks.org/wiki/High_School_Chemistry/Using_Data but i don't understand it still can i have step by step help

OpenStudy (micahm):

@ganeshie8

OpenStudy (micahm):

@confluxepic

OpenStudy (micahm):

@Frostbite

OpenStudy (micahm):

@Coolsector

OpenStudy (micahm):

@sleepyjess

OpenStudy (micahm):

@sleepyjess

OpenStudy (frostbite):

What experiment is this?

OpenStudy (micahm):

can i send you the doc i don't think it's an experiment

OpenStudy (frostbite):

Well no need to. Have you tried plotting the data?

OpenStudy (micahm):

they show a graph

OpenStudy (frostbite):

Might be a good idea to show the document (pdf please) or you use the draw tool. Give me all the information you got. Basically what it sounds like to me is a P vs T graph for a gas in a closed container. To which we can approximate with the ideal gas equation: \[\Large pV=nRT \rightarrow p=\frac{ nR }{ V } \times T\] Assuming that you get a somewhat linear graph we can use the ideal gas equation for fitting as the ideal gas takes the linear form above with the slope: \[\Large \frac{ dp }{ dT }=\frac{ nR }{V }\] That is the theory part for it. So by taking the data and putting into a graph we could get something like this yes:  Doing a linear fit we could get